OpenStudy (anonymous):

solve by the elimination method 0.05x+0.25y=11 0.15x+0.05y=12

OpenStudy (anonymous):

0.05x+0.25y=11---(a) 0.15x+0.05y=12---(b) multiply equation (a) with 3 u get 0.15x+0.75y=33---(c) subtract eqn (b) from (c) 0.15x+0.75y=33---(c) (-) 0.15x+0.05y=12---(b) ---------------------- 0 + 0.70y = 21 0.70y =21 y=21-0.70 = 30 put y= 30 in (a) 0.05x+0.25(30)=11 0.05x=11-7.5=3.5 0.05x=3.5 x=70 so x=70;y=30

OpenStudy (anonymous):

that solution is not correct. Is this in simplist form?